# Amortized Inference of Causal Models via Conditional Fixed-Point Iterations

### v3

@@ Abstract
Structural Causal Models (SCMs) offer a principled framework to reason about interventions and support out-of-distribution generalization, which are key goals in scientific discovery. However, the task of learning SCMs from observed data poses formidable challenges, and often requires training a separate model for each dataset. In this work, we propose amortized inference of SCMs by training a single model on multiple datasets sampled from different SCMs. We first use a transformer-based architecture for amortized learning of dataset embeddings, and then extend the Fixed-Point Approach (FiP) (Scetbon et al., 2024) to infer SCMs conditionally on their dataset embeddings. As a byproduct, our method can generate observational and interventional data from novel SCMs at inference time, without updating parameters. Empirical results show that our amortized procedure performs on par with baselines trained specifically for each dataset on both in and out-of-distribution problems, and also outperforms them in scare data regimes. Equal Contribution. Correspondence to: meyerscetbon@gmail.com. This work was done when Divyat Mahajan was an intern at Microsoft Research. Cheng Zhang worked on this project when she was affiliated with Microsoft Research and she is currently with Meta.
@@ 1 Introduction
Despite these advances, a major limitation remains: each new dataset requires training a specific model, that prevents the transfer of causal knowledge across datasets. Amortized inference offers a solution by learning a single model that can generalize across instances of the same optimization problem by exploiting their shared structure (Andrychowicz et al., 2016; Gordon et al., 2019). This results in models that can quickly adapt to new instances at test time (Finn et al., 2017). Amortized inference has been shown success in several challenging tasks, like bayesian posterior estimation (Garnelo et al., 2018; Müller et al., 2021), sampling from unnormalized densities (Akhound-Sadegh et al., 2024; Sendera et al., 2024), as well as causal structure learning (Lorch et al., 2022; Ke et al., 2022), which is more aligned with our paper.
In this work, we tackle the novel problem of amortized inference of causal mechanisms for additive noise SCMs. We propose a two-step approach where we first learn dataset embeddings via in-context learning (Garg et al., 2022) to represent the task-specific information. These embeddings are then used to condition the fixed-point (FiP) approach (Scetbon et al., 2024) for modeling causal mechanisms. This conditional modification, termed Cond-FiP, enables the model to adapt the causal mechanism for each specific instance (Figure 1). Our key contributions are highlighted below.
@@
  • We show empirically [^1] The code is available on Github: microsoft/causica. that Cond-FiP achieves similar performances as the state-of-the-art (SOTA) approaches trained from scratch for each dataset on both in and out-of-distribution (OOD) problems. Further, Cond-FiP obtains better results than baselines in scare data regimes, due to its amortized inference procedure.
@@ 2.1 Brief Overview of Amortized Inference
Amortized inference aims to learn a shared inference mechanism across multiple tasks that enables fast adaptation to new tasks at test time. Consider a task $T$ that defines a distribution over inputs (${\bm{Z}}$) and targets (${\bm{Y}}$), i.e, ${\bm{Z}},{\bm{Y}}\sim{\mathbb{P}}_{T}$. Given a collection of tasks $\big{(}T^{(k)}\big{)}_{k=1}^{K}$ and some objective function $L$, the goal is to learn a model ${\mathcal{T}}_{\theta}$ shared across tasks as follows:
@@
where ${\bm{I}}^{(k)}$ denotes additional context for task $T^{(k)}$, such as dataset with samples $[{\bm{Z}}_{1},\cdots,{\bm{Z}}_{n}]$. Instead of retraining from scratch, the model should leverage the context ${\bm{I}}^{{}^{\prime}}$ to adapt to the task $T^{{}^{\prime}}$.
A classic approach for this is meta-learning (Andrychowicz et al., 2016; Finn et al., 2017), that utilizes context ${\bm{I}}^{{}^{\prime}}$ by task-specific finetuning. These methods typically learn a shared initialization that is refined for a specific task via few gradient steps in an inner optimization loop.
In contrast, in-context learning (ICL) (Müller et al., 2021; Xie et al., 2021; Garg et al., 2022) avoids this inner loop by using transformer-based architectures. By attending to the context ${\bm{I}}^{{}^{\prime}}$ during the forward pass, ICL methods adapt to a specific task without any parameter updates. This ability arises from the observation that transformers can implicitly approximate learning algorithms such as gradient descent within their activation dynamics (Akyürek et al., 2022; Von Oswald et al., 2023).
@@ 2.2 Problem Setup
Hence, an SCM ${\mathcal{S}}({\mathbb{P}}_{{\bm{N}}},{\bm{\mathcal{G}}},{\bm{F}})$ describes the data-generation process of ${\bm{X}}:=[X_{1},\cdots,X_{d}]\sim{\mathbb{P}}_{{\bm{X}}}$ from the noise variables ${\bm{N}}:=[N_{1},\cdots,N_{d}]\sim{\mathbb{P}}_{{\bm{N}}}$ via the function ${\bm{F}}:=[F_{1},\cdots,F_{d}]$, and a graph ${\bm{\mathcal{G}}}\in\{0,1\}^{d\times d}$ indicating the parents of each $X_{i}$, that is $[{\bm{\mathcal{G}}}]_{i,j}:=1$ if $X_{j}\in\emph{PA}(X_{i})$. We make the following assumptions about SCMs.
@@
While the first assumption is pretty standard, we make the ANM assumption for training the proposed dataset encoder in Section 3.1.
@@ 3.1 Dataset Encoder
The objective of this section is to develop a method capable of producing efficient latent representations of datasets. To achieve this, we propose to train an encoder that predicts the noise samples from their associated observations given the causal structures via in-context learning.
@@
Encoder Architecture. Following (Lorch et al., 2021; Scetbon et al., 2024), we encode datasets using a transformer-based architecture that alternates attention over both sample and node dimension. Given a dataset $D_{\bm{X}}$, we first apply a linear embedding $L(D_{\bm{X}})\in\mathbb{R}^{n\times d\times d_{h}}$, where $d_{h}$ is the hidden dimension. The encoder $E$ then applies transformer blocks, each comprising self-attention followed by an MLP  (Vaswani et al., 2017), where the attention mechanism is applied either across the samples $n$ or the nodes $d$ alternately. Recall the standard self-attention is defined as
@@
Training Procedure. We minimize the mean squared error (MSE) of predicting the target $D_{\bm{N}}$ from the input ($D_{\bm{X}},{\bm{\mathcal{G}}}$) over the distribution of SCMs ${\mathbb{P}}_{{\mathcal{S}}}$ available during training:
${\mathbb{E}}_{{\mathcal{S}}\sim{\mathbb{P}}_{{\mathcal{S}}}}||D_{\bm{N}}-H\circ E(L(D_{\bm{X}}),{\bm{\mathcal{G}}})||^{2}_{2}\;.$
@@
${\mathbb{E}}_{{\mathcal{S}}\sim{\mathbb{P}}_{{\mathcal{S}}}}||\bm{F}(D_{\bm{X}})-H\circ E(L(D_{\bm{X}}),{\bm{\mathcal{G}}})||^{2}_{2}\;.$
Note that ANM assumption provides a simplified true mapping from data to noise as $x\to x-F(x)$, which is difficult to obtain in general SCMs. Please check Section A.2 for more details on justification for ANMs and why recovering noise is equivalent to learning the inverse SCM.
Inference. Given a new dataset $D_{\bm{X}}$ and its causal graph ${\bm{\mathcal{G}}}$, encoder provides us with the dataset embedding $\mu(D_{\bm{X}},{\bm{\mathcal{G}}}):=E(L(D_{\bm{X}}),{\bm{\mathcal{G}}})\in\mathbb{R}^{n\times d\times d_{h}}$.
@@ 3.2 Cond-FiP: Conditional Fixed-Point Decoder
We now present the modification of FiP that uses the learned dataset embeddings $\mu(D_{\bm{X}},{\bm{\mathcal{G}}})$ for amortized inference of causal mechanisms.
Training Setting. Analogous to the encoder training setup, we assume that we have access to a distribution of SCMs ${\mathcal{S}}({\mathbb{P}}_{{\bm{N}}},{\bm{\mathcal{G}}},{\bm{F}})\sim{\mathbb{P}}_{{\mathcal{S}}}$ at training time, from which we can extract empirical representations $(D_{\bm{X}},{\bm{\mathcal{G}}})$. Our goal is to train ${\mathcal{T}}$ such that given the context $(D_{\bm{X}},{\bm{\mathcal{G}}})$ from an SCM ${\mathcal{S}}({\mathbb{P}}_{{\bm{N}}},{\bm{\mathcal{G}}},{\bm{F}})\sim{\mathbb{P}}_{{\mathcal{S}}}$, the induced conditional function $\bm{z}\in\mathbb{R}^{d}\to\mathcal{T}(\bm{z},D_{\bm{X}},{\bm{\mathcal{G}}})\in\mathbb{R}^{d}$ approximates the true causal mechanisms $\bm{F}:\bm{z}\in\mathbb{R}^{d}\to\bm{F}(\bm{z})\in\mathbb{R}^{d}$ (E.q. 3).
Decoder Architecture. The design of our decoder is based on the FiP architecture for fixed-point SCM learning, with two major differences: (1) we use the dataset embeddings $\mu(D_{\bm{X}},{\bm{\mathcal{G}}})$ as a high dimensional codebook to embed the nodes, and (2) we leverage adaptive layer norm operators (Peebles and Xie, 2023) in the transformer blocks of FiP to enable conditional attention mechanisms.
Conditional Embedding. The key change of our decoder compared to the original FiP is in the embedding of the input. FiP proposes to embed a data point $\bm{z}:=[z_{1},\dots,z_{d}]\in\mathbb{R}^{d}$ into a high dimensional space using a learnable codebook $\bm{C}:=[C_{1},\dots,C_{d}]^{T}\in\mathbb{R}^{d\times d_{h}}$ and positional embedding $\bm{P}:=[P_{1},\dots,P_{d}]^{T}\in\mathbb{R}^{d\times d_{h}}$, from which they define:
@@
$\bm{C}(D_{\bm{X}},{\bm{\mathcal{G}}}):=\mu(D_{\bm{X}},{\bm{\mathcal{G}}})W_{\bm{C}}$
$\bm{P}(D_{\bm{X}},{\bm{\mathcal{G}}}):=\mu(D_{\bm{X}},{\bm{\mathcal{G}}})W_{\bm{P}}$
@@
$\bm{z}_{\text{emb}}:=$ $[z_{1}*C_{1}(D_{\bm{X}},{\bm{\mathcal{G}}}),\dots,z_{d}*C_{d}(D_{\bm{X}},{\bm{\mathcal{G}}})]^{T}+\bm{P}(D_{\bm{X}},{\bm{\mathcal{G}}})\in\mathbb{R}^{d\times d_{h}}$
Adaptive Transfomer Block. Once an input $\bm{z}\in\mathbb{R}^{d}$ has been embedded as $\bm{z}_{\text{emb}}\in\mathbb{R}^{d\times d_{h}}$, FiP models SCMs by simulating the reconstruction of the data from noise. Starting from $\bm{n}_{0}\in\mathbb{R}^{d\times d_{h}}$ a learnable parameter, they propose to update the current noise $L\geq 1$ times by computing:
@@
where $h$ refers to the MLP block, and for clarity, we omit both the layer’s dependence on its parameters and the inclusion of layer normalization in the notation. Note that here FiP considers the DAG-Attention mechanism (details in Section A.1) in order to correctly model the root nodes of the SCM. To obtain a conditional formulation, we first replace the starting noise $\bm{n}_{0}$ with $\bm{n}_{0}:=\mu(D_{\bm{X}},{\bm{\mathcal{G}}})W_{\bm{n}_{0}}\in\mathbb{R}^{d\times d_{h}}$, where $W_{\bm{n}_{0}}\in\mathbb{R}^{d_{h}\times d_{h}}$ is a learnable parameter. Then we add adaptive layer normalization operators (Peebles and Xie, 2023) to both attention and MLP blocks, where each scale or shift is obtained by applying a 1 hidden-layer MLP to the embedding $\mu(D_{\bm{X}},{\bm{\mathcal{G}}})$.
Projection. To project back the latent representation of $\bm{z}$ obtained from previous stages, $\bm{n}_{L}\in\mathbb{R}^{d\times d_{h}}$, we use a linear operation to get $\widehat{\bm{z}}=\bm{n}_{L}W_{\text{out}}\in\mathbb{R}^{d}$, where $W_{\text{out}}\in\mathbb{R}^{d_{h}}$ is learnable.
Training Procedure. The result of forward pass can be summarized as $\widehat{\bm{z}}={\mathcal{T}}(\bm{z},D_{\bm{X}},{\bm{\mathcal{G}}})$, where we omit the dependence of $\widehat{\bm{z}}$ on context $(D_{\bm{X}},{\bm{\mathcal{G}}})$ for simplicity. We train the model ${\mathcal{T}}$ by minimizing the reconstruction error of the true causal mechanisms estimated by our model over the distribution of SCMs ${\mathbb{P}}_{{\mathcal{S}}}$, as shown below.

### v4

@@ Abstract
Structural Causal Models (SCMs) offer a principled framework to reason about interventions and support out-of-distribution generalization, which are key goals in scientific discovery. However, the task of learning SCMs from observed data poses formidable challenges, and often requires training a separate model for each dataset. In this work, we propose an amortized inference framework that trains a single model to predict the causal mechanisms of SCMs conditioned on their observational data and causal graph. We first use a transformer-based architecture for amortized learning of dataset embeddings, and then extend the Fixed-Point Approach (FiP) to infer the causal mechanisms conditionally on their dataset embeddings. As a byproduct, our method can generate observational and interventional data from novel SCMs at inference time, without updating parameters. Empirical results show that our amortized procedure performs on par with baselines trained specifically for each dataset on both in and out-of-distribution problems, and also outperforms them in scarce data regimesEqual Contribution. Correspondence to: divyatmahajan@gmail.com.This work was done when DM was an intern at Microsoft Research. JG, AH, CZ, and MS worked on this project when they were affiliated with Microsoft Research. .
@@ 1 Introduction
Despite these advances, a major limitation remains: each new dataset requires training a specific model, that prevents the transfer of causal knowledge across datasets. Amortized inference offers a solution by learning a single model that can generalize across instances of the same optimization problem by exploiting their shared structure (Andrychowicz et al., 2016; Gordon et al., 2019). This results in models that can quickly adapt to new instances at test time (Finn et al., 2017). Amortized inference has shown success in several challenging tasks, like bayesian posterior estimation (Garnelo et al., 2018; Müller et al., 2021), sampling from unnormalized densities (Akhound-Sadegh et al., 2024; Sendera et al., 2024), as well as causal structure learning (Lorch et al., 2022; Ke et al., 2022), which is more aligned with our paper.
In this work, we tackle the novel problem of amortized inference of causal mechanisms for additive noise SCMs. While prior research has primarily focused on amortized approaches for causal discovery (Lorch et al., 2022; Dhir et al., 2024) or treatment effect estimation (Nilforoshan et al., 2023; Zhang et al., 2023), our goal instead is to train a single model capable of inferring the causal mechanisms of novel SCMs, given their observational data and associated causal graph. We propose a two-step approach where we first learn dataset embeddings via in-context learning (Garg et al., 2022) to represent the task-specific information. These embeddings are then used to condition the fixed-point (FiP) approach (Scetbon et al., 2024) for modeling causal mechanisms. This conditional modification, termed Cond-FiP, enables the model to adapt the causal mechanism for each specific instance (Figure˜1). Our key contributions are highlighted below.
@@
  • We show empirically  [^1] The code is available on Github: microsoft/causica. that Cond-FiP achieves similar performances as the state-of-the-art (SOTA) approaches trained from scratch for each dataset on both in and out-of-distribution (OOD) problems. Further, Cond-FiP obtains better results than baselines in scarce data regimes, due to its amortized inference procedure.
1.1 Related Works
Amortized Causal Learning.
Amortized inference has gained traction in causality research, particularly for structure learning. Early works by Lorch et al. (2022) and Ke et al. (2022) introduced transformer-based models trained on multiple synthetic datasets using supervised objectives for amortized inference of causal structure. Their approach aligns with recent works on in-context learning of function classes via transformers (Müller et al., 2021; Akyürek et al., 2022; Garg et al., 2022; Von Oswald et al., 2023). Subsequent improvements targeted OOD generalization (Wu et al., 2024), bayesian causal discovery (Dhir et al., 2024), and applications to gene regulatory networks (Ke et al., 2023; Kim et al., 2025). Beyond structure learning, amortized methods have been developed for ATE/CATE estimation (Nilforoshan et al., 2023; Zhang et al., 2023; Sauter et al., 2025; Bynum et al., 2025; Balazadeh et al., 2025; Robertson et al., 2025; Ma et al., 2025; Dhir et al., 2025), model selection (Gupta et al., 2023), and partial causal discovery tasks such as learning topological order (Scetbon et al., 2024). However, amortized inference of causal mechanisms in SCMs remains unaddressed, which is the central focus of our work.
Autoregressive Causal Learning. Most causal discovery methods focus first on structure learning (Chickering, 2002; Peters et al., 2014; Zheng et al., 2018), followed by per-node maximum likelihood estimation to recover the causal mechanisms (Blöbaum et al., 2022). In constrast, recent works on causal autoregressive flows (Khemakhem et al., 2021; Geffner et al., 2022; Javaloy et al., 2023) focus on SOTA normalizing flow based generative models to infer causal mechanisms. Further, FiP (Scetbon et al., 2024) modeled SCMs as fixed-point problems over causal (topological) ordering of nodes using transformer-based architectures. These approaches efficiently learn SCMs but require training a separate model per dataset. In this work, we remove this limitation by extending FiP to enable amortized inference of causal mechanisms across different SCMs.
@@ 2.1 Brief Overview of Amortized Inference
Amortized inference aims to learn a shared inference mechanism across multiple tasks that enables fast adaptation to new tasks at test time. We motivate amortized inference through the following example, consider the task of predicting the motion of objects on different planets. While trajectories vary across planets across due to differences in gravitational constants, the underlying physical laws remain the same. Rather than training a new model from scratch for each planet, we can exploit this shared structure to rapidly adapt our predictions to new settings. This is the core idea behind amortized inference, leveraging common patterns across training tasks to enable fast and efficient adaptation to novel ones.
Consider a task $T$ that defines a distribution over inputs ($\bm{Z}$}) and targets ($\bm{Y}$}); ${\bm{Z}},{\bm{Y}}\sim{\mathbb{P}}_{T}$. Given a collection of tasks $\big(T^{(k)}\big)_{k=1}^{K}$ and some objective function $L$, the goal is to learn a shared model ${\mathcal{T}}_{\theta}$ across tasks:
@@
where ${\bm{I}}^{(k)}$ denotes additional context for task $T^{(k)}$, such as dataset with samples ${\bm{Z}}_{1},\cdots,{\bm{Z}}_{n}]$. Instead of retraining from scratch, the model should leverage the context $\bm{I}$$^{{}^{\prime}}$ to adapt to the task $T^{{}^{\prime}}$.
A classic approach towards this is meta-learning (Andrychowicz et al., 2016; Finn et al., 2017; Hospedales et al., 2021), which leverages the context $\bm{I}$$^{{}^{\prime}}$ by task-specific finetuning. These methods typically learn a shared initialization that is adapted for a new task via few gradient steps in an inner optimization loop. In contrast, in-context learning (ICL) (Müller et al., 2021; Xie et al., 2021; Garg et al., 2022; Akyürek et al., 2024) avoids this inner loop by using transformer-based architectures. By attending to the context $\bm{I}$$^{{}^{\prime}}$ during the forward pass, ICL methods adapt to a new task without any parameter updates. This capability is often attributed to transformers’ ability to implicitly approximate learning algorithms such as gradient descent within their activations (Akyürek et al., 2022; Von Oswald et al., 2023).
While ICL is often used to describe "emergent" test-time adaptation in large language models (Brown et al., 2020), where the training objective does not explicitly involve those tasks. Here, we focus on its formulation as prior-fitted networks (PFNs) (Hollmann et al., 2022; Robertson et al., 2024), where transformers are pretrained on synthetic datasets generated from a simulator that implicitly defines a prior over tasks.
@@ 2.2 Problem Setup
Hence, an SCM ${\mathcal{S}}({\mathbb{P}}_{{\bm{N}}},{\bm{\mathcal{G}}},{\bm{F}})$ describes the data-generation process of ${\bm{X}}:=[X_{1},\cdots,X_{d}]\sim{\mathbb{P}}_{{\bm{X}}}$ from the noise variables ${\bm{N}}:=[N_{1},\cdots,N_{d}]\sim{\mathbb{P}}_{{\bm{N}}}$ via the function ${\bm{F}}:=[F_{1},\cdots,F_{d}]$, and a graph ${\bm{\mathcal{G}}}\in\{0,1\}^{d\times d}$ indicating the parents of each $X_{i}$, that is ${\bm{\mathcal{G}}}]_{i,j}:=1$ if $X_{j}\in\emph{PA}(X_{i})$. We make the following assumptions about SCMs.
@@
While the first assumption is pretty standard, we make the ANM assumption for training the proposed dataset encoder in Section˜3.1.
@@ 3.1 Dataset Encoder
The objective of this section is to develop a method capable of producing efficient latent representations of datasets. To achieve this, we propose to train an encoder that predicts the noise samples from their associated observations given the causal structures via in-context learning (pseudo code in Algorithm 1Section˜A.4).
@@
Encoder Architecture.
Following (Lorch et al., 2021; Scetbon et al., 2024), we encode datasets using a transformer-based architecture that alternates attention over both sample and node dimension. Given a dataset $D_{\bm{X}}$, we first apply a linear embedding $L(D_{\bm{X}})\in\mathbb{R}^{n\times d\times d_{h}}$, where $d_{h}$ is the hidden dimension. The encoder $E$ then applies transformer blocks, each comprising self-attention followed by an MLP  (Vaswani et al., 2017), where the attention mechanism is applied either across the samples $n$ or the nodes $d$ alternately. Recall the standard self-attention is defined as
@@
Training Procedure.
We minimize the mean squared error (MSE) of predicting the target $D_{\bm{N}}$ from the input ($D_{\bm{X}},{\bm{\mathcal{G}}}$) over the distribution of SCMs ${\mathbb{P}}_{{\mathcal{S}}}$ available during training:
[…]
@@
[…]
Note that ANM assumption provides a simplified true mapping from data to noise as $x\to x-F(x)$, which is difficult to obtain in general SCMs. Please check Section˜A.2 for more details on justification for ANMs and why recovering noise is equivalent to learning the inverse SCM.
Inference.
Given a new dataset $D_{\bm{X}}$ and its causal graph $\bm{\mathcal{G}}$}, encoder provides us with the dataset embedding $\mu(D_{\bm{X}},{\bm{\mathcal{G}}}):=E(L(D_{\bm{X}}),{\bm{\mathcal{G}}})\in\mathbb{R}^{n\times d\times d_{h}}$.
@@ 3.2 Cond-FiP: Conditional Fixed-Point Decoder
We now present the modification of FiP that uses the learned dataset embeddings $\mu(D_{\bm{X}},{\bm{\mathcal{G}}})$ for amortized inference of causal mechanisms (pseudo code in Algorithm˜2Section˜A.4).
Training Setting.
Analogous to the encoder training setup, we assume that we have access to a distribution of SCMs ${\mathcal{S}}({\mathbb{P}}_{{\bm{N}}},{\bm{\mathcal{G}}},{\bm{F}})\sim{\mathbb{P}}_{{\mathcal{S}}}$ at training time, from which we can extract empirical representations $(D_{\bm{X}},{\bm{\mathcal{G}}})$. Our goal is to train $\mathcal{T}$} such that given the context $(D_{\bm{X}},{\bm{\mathcal{G}}})$ from an SCM ${\mathcal{S}}({\mathbb{P}}_{{\bm{N}}},{\bm{\mathcal{G}}},{\bm{F}})\sim{\mathbb{P}}_{{\mathcal{S}}}$, the induced conditional function $\bm{z}\in\mathbb{R}^{d}\to\mathcal{T}(\bm{z},D_{\bm{X}},{\bm{\mathcal{G}}})\in\mathbb{R}^{d}$ approximates the true causal mechanisms $\bm{F}:\bm{z}\in\mathbb{R}^{d}\to\bm{F}(\bm{z})\in\mathbb{R}^{d}$ (E.q. 3).
Decoder Architecture.
The design of our decoder is based on the FiP architecture for fixed-point SCM learning, with two major differences: (1) we use the dataset embeddings $\mu(D_{\bm{X}},{\bm{\mathcal{G}}})$ as a high dimensional codebook to embed the nodes, and (2) we leverage adaptive layer norm operators (Peebles & Xie, 2023) in the transformer blocks of FiP to enable conditional attention mechanisms.
Conditional Embedding.
The key change of our decoder compared to the original FiP is in the embedding of the input. FiP proposes to embed a data point $\bm{z}:=[z_{1},\dots,z_{d}]\in\mathbb{R}^{d}$ into a high dimensional space using a learnable codebook $\bm{C}:=[C_{1},\dots,C_{d}]^{T}\in\mathbb{R}^{d\times d_{h}}$ and positional embedding $\bm{P}:=[P_{1},\dots,P_{d}]^{T}\in\mathbb{R}^{d\times d_{h}}$, from which they define:
@@
$\bm{C}(D_{\bm{X}},{\bm{\mathcal{G}}})=\mu(D_{\bm{X}},{\bm{\mathcal{G}}})W_{\bm{C}}$
$\bm{P}(D_{\bm{X}},{\bm{\mathcal{G}}})=\mu(D_{\bm{X}},{\bm{\mathcal{G}}})W_{\bm{P}}$
@@
$\bm{z}_{\text{emb}}$= $[z_{1}*C_{1}(D_{\bm{X}},{\bm{\mathcal{G}}}),\dots,z_{d}*C_{d}(D_{\bm{X}},{\bm{\mathcal{G}}})]^{T}+\bm{P}(D_{\bm{X}},{\bm{\mathcal{G}}})\in\mathbb{R}^{d\times d_{h}}$
Adaptive Transfomer Block.
Once an input $\bm{z}\in\mathbb{R}^{d}$ has been embedded as $\bm{z}_{\text{emb}}\in\mathbb{R}^{d\times d_{h}}$, FiP models SCMs by simulating the reconstruction of the data from noise. Starting from $\bm{n}_{0}\in\mathbb{R}^{d\times d_{h}}$ a learnable parameter, they propose to update the current noise $L\geq 1$ times by computing:
@@
where $h$ refers to the MLP block, and for clarity, we omit both the layer’s dependence on its parameters and the inclusion of layer normalization in the notation. Note that here FiP considers the DAG-Attention mechanism (details in Section˜A.1) in order to correctly model the root nodes of the SCM. To obtain a conditional formulation, we first replace the starting noise $\bm{n}_{0}$ with $\bm{n}_{0}:=\mu(D_{\bm{X}},{\bm{\mathcal{G}}})W_{\bm{n}_{0}}\in\mathbb{R}^{d\times d_{h}}$, where $W_{\bm{n}_{0}}\in\mathbb{R}^{d_{h}\times d_{h}}$ is a learnable parameter. Then we add adaptive layer normalization operators (Peebles & Xie, 2023) to both attention and MLP blocks, where each scale or shift is obtained by applying a 1 hidden-layer MLP to the embedding $\mu(D_{\bm{X}},{\bm{\mathcal{G}}})$.
Projection.
To project back the latent representation of $\bm{z}$ obtained from previous stages, $\bm{n}_{L}\in\mathbb{R}^{d\times d_{h}}$, we use a linear operation to get $\widehat{\bm{z}}=\bm{n}_{L}W_{\text{out}}\in\mathbb{R}^{d}$, where $W_{\text{out}}\in\mathbb{R}^{d_{h}}$ is learnable.
Training Procedure.
The result of forward pass can be summarized as $\widehat{\bm{z}}={\mathcal{T}}(\bm{z},D_{\bm{X}},{\bm{\mathcal{G}}})$, where we omit the dependence of $\widehat{\bm{z}}$ on context $(D_{\bm{X}},{\bm{\mathcal{G}}})$ for simplicity. We train the model $\mathcal{T}$} by minimizing the reconstruction error of the true causal mechanisms estimated by our model over the distribution of SCMs ${\mathbb{P}}_{{\mathcal{S}}}$, as shown below.
@@
Remark on ANM assumption.
Though our method relies on the ANM assumption for encoder training (Appendix A.2), we do not need this assumption for decoder training! An interesting direction for future work is to develop more general encoder training strategies, such as using self-supervised learning for dataset encoding. Another option is to pursue end-to-end training of Cond-FiP, or adopt curriculum learning by first training the encoder to predict noise variables (as a "pretraining" step), and then fine-tuning it using the decoder’s reconstruction loss on more realistic SCMs. However, we consider these extensions beyond the scope of the current work.